OpenStudy (anonymous):

A recipe that makes 7 servings calls for 1/2 cup of apple juice and 2/3 cup of cranberry juice. How many cups of juice are there per serving?

OpenStudy (anonymous):

2/3 + 1/2 3/6 + 4/6 = 7/6 7/6 divided by 7 = 0.1666666666666667

OpenStudy (anonymous):

approx .16 cups of juice/serving

OpenStudy (anonymous):

Is that 1/6? :-)

OpenStudy (anonymous):

yea.... sorry I do better with decimals so I put my numbers in decimals...but 1/6 is the same thing

OpenStudy (anonymous):

Ok :-) Thank you soooo much ily